OpenStudy (anonymous):

WHO EVER ANSWERS THIS ILL GIVE YOU A MEDAL Using the equation y equals x divided by five, what is the value of y when x is twenty? y = x over 5 x y 5 1 15 3 20 ? 40 8 4 5 6 7

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

answer is 4?

OpenStudy (anonymous):

If Y=x/5 then 5y=x so if 4=20/5 then 4*5=20. The answer is 4.

OpenStudy (anonymous):

Ok thank you so much.

OpenStudy (anonymous):

if x=20 than 20 is devided by 5 than it equals 4 because 4x5=20

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

id choose you both but I cant lol

OpenStudy (anonymous):

I was beat

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

@lakotareid1232 anyway I think it doesn't have importance for you because you've got enough points :D @kopster123 bye :)

OpenStudy (anonymous):

that's rude bro